# https://www.dekelgalor.com/noise2image

### v1

@@
Abstract
Event cameras capture changes of intensity over time as a stream of ‘events’ and generally cannot measure intensity itself; hence, they are only used for imaging dynamic scenes. However, fluctuations due to random photon arrival inevitably trigger noise events, even for static scenes. While previous efforts have been focused on filtering out these undesirable noise events to improve signal quality, we find that, in the photon-noise regime, these noise events are correlated with the static scene intensity. We analyze the noise event generation and model its relationship to illuminance. Based on this understanding, we propose a method, called Noise2Image, to leverage the illuminance-dependent noise characteristics to recover the static parts of a scene, which are otherwise invisible to event cameras. We experimentally collect a dataset of noise events on static scenes to train and validate Noise2Image. Our results show that Noise2Image can robustly recover intensity images solely from noise events, providing a novel approach for capturing static scenes in event cameras, without additional hardware.
Keywords:
dynamic vision sensor, event-based camera, image reconstruction, noise characterization
@@ 1 Introduction
Event cameras [34], also known as neuromorphic cameras or dynamic vision sensors, are an emerging modality for capturing dynamic scenes. Their ability to capture data at a much faster rate than conventional cameras have led to their application in high-speed navigation, augmented reality, and real-time 3D reconstruction [19]. Unlike a conventional CMOS camera that outputs intensity images at fixed intervals, an event camera detects brightness changes at each pixel asynchronously. When the brightness change at a pixel exceeds some threshold, an event is recorded. The output from an event consists of three elements: a timestamp, the spatial coordinate of the triggered pixel, and a binary polarity, indicating whether it is an increase or decrease of brightness.
@@
While this scheme allows event cameras to operate beyond conventional framerates, it makes them blind to the stationary components of a scene, which induce no brightness changes over time. This issue is especially prevalent when the camera is not moving and therefore provides no information about the static background. Even though event cameras are not designed to capture an intensity image, it is often needed for downstream applications such as initializing motion tracking algorithms [2]. To mitigate this issue, some event cameras include a conventional frame-based sensor in the pixel circuit to simultaneously image both events and traditional intensity images [9, 7]. Alternatively, when circuit-level modification of the event camera is not possible, a frame-based camera can be installed in parallel, using either a beam splitter [24, 58, 57] or additional view registration [51]. Both of these solutions introduce additional hardware, increasing cost, complexity, size, and power consumption.
Figure 1: Schematic of the Noise2Image pipeline. Recorded events are first separated into noise events and signal events using an existing event denoiser. Signal events are triggered by intensity changes of the scene, which can be fed into existing event-to-video reconstruction methods. The noise events are then used to reconstruct the static scene intensity with our Noise2Image method. This relies on characterizing the relationship between noise events and using learned priors to resolve ambiguities. Data in this figure was captured in an outdoor environment late afternoon.
@@
Our work leverages the fact that, even when the scene is static, event cameras still produce noise events. We focus on low and moderate-brightness regimes (e.g. room light, outdoor sunset) [23, 22, 21] in which the dominant source of noise is photon noise — random fluctuations in the photon arrival process. In contrast, the high-brightness regime (e.g. outdoor daylight) includes leakage noise events [36], which we do not model here. While the photon noise is well-studied in the context of frame-based scenes, those events triggered by photon noise are commonly deemed as part of the general background noise activity of the event-based sensor, to be filtered out.
@@
In this work we propose a method called Noise2Image that can reconstruct a static scene from its event noise statistics, with no hardware modifications and negligible computational overhead. First, we derive a statistical noise model describing how noise event generation correlates with scene intensity, which shows a good correspondence with our experimental measurements. Unlike in conventional sensors, where photon noise grows with the signal, we find that for event cameras, the number of events triggered by photon noise is mostly negatively correlated with the illuminance level due to the logarithmic sensitivity of the sensor. Imaging the static scene then amounts to inverting this intensity-to-noise process. However, the mapping is one-to-many, so not directly invertible; thus, we rely on a learned prior to resolve ambiguities. To train and validate our method, we experimentally collect a dataset of event recordings on static scenes. We demonstrate that Noise2Image can recover photo-realistic images from noise events alone, or can be used to recover the static parts of scenes with dynamics. In addition to testing on in-distribution data, we demonstrate the robustness of this approach with out-of-distribution testing data and live scenes (Fig. 1).
Our contributions are summarized as follows:
  • We characterize noise event generation due to photon noise and derive a mathematical model describing the statistical relationship between noise events and pixel illuminance.
  • We propose the Noise2Image method to recover the intensity image of a static scene from a recording of noise events using a learned prior.
  • We collect a noise-events-to-image (NE2I) dataset with recordings of noise events paired with the corresponding intensity images to train and validate our method.
  • We show that Noise2Image is complementary to event-to-video reconstruction methods (E2VID), enabling recovery of both static and dynamic parts of a scene.
@@ 2 Related Work
2.0.1 Event-to-video reconstruction.
Event-to-video reconstruction (E2VID) is a class of methods that recover high frame-rate video from event recordings. These methods can capture high-speed dynamics without the motion blur that plagues traditional frame-based cameras. Because event cameras only capture scene changes, many E2VID approaches include one or two traditional frame-based images, using events to deblur [43, 27, 39, 49], synthesize adjacent frames [50], or interpolate temporally [48, 47, 55, 52].
In contrast, a more challenging class of E2VID reconstruction only uses events for video reconstruction. Because the initial scene intensity is unknown and only the relative intensity changes are measured, reconstruction requires either explicit modeling of spatiotemporal relationships [4] or deep neural networks as a data prior to fill in the missing information [42, 46, 44, 53, 10, 40, 17]. Because it is difficult to collect event recordings paired with the corresponding frame-based videos at scale, the data prior is obtained through synthetic data generation [42, 46] using an event camera simulator [41]. While E2VID methods are effective for recordings with object motion or camera motion, the reconstruction of static scenes or regions is still out of reach since no events are thought to be triggered without motion. Our Noise2Image method can be considered complementary to E2VID; one might use Noise2Image to recover the static parts of the scene and E2VID to recover the dynamic parts.
2.0.2 Event camera noise characterization.
An event recording often contains a number of events not associated with intensity changes, which are termed noise events or background activity [12, 22]. Noise events are attributed to two main sources: photon noise and leakage current [36, 21, 23]. Photon noise is the dominant source of noise in low-brightness conditions [31, 21], while leakage noise events dominates in high-brightness conditions [36, 21]. Although the noise model of event cameras is less studied than traditional sensors, in the lower-light regime, it is generally believed that noise events become less likely to trigger as intensity increases [14, 26]. In event camera simulators, the events triggered by photon noise are modeled as a Poisson process, with the noise event rate linearly decreasing with intensity [26]. The intensity dependency of noise events has been experimentally measured in [21], and shows a non-monotonic relationship with intensity in low light. Our experimental results are consistent with this trend, and we derive a theoretical noise model which explains the relationship between intensity and noise events.
2.0.3 Event denoising.
@@
2.0.4 Event camera datasets.
@@ 3 Modeling Noise Event Statistics
where $t_{0}$ is the timestamp for the most recent event at the same pixel. The logarithmic sensitivity ensures that this triggering condition is adaptive to the brightness level, e.g. a tenfold increase from 1 lux to 10 lux would result in a similar response as a tenfold increase from 100 lux to 1000 lux. As pixels operate asynchronously, we will proceed by considering events at an arbitrary pixel and drop the $\left(x,y\right)$ notation for simplicity.
@@
To trigger a positive event, we need $\log\left(n\right)-\log\left(n_{0}\right)>\epsilon$, which can be re-written as $n-n_{0}e^{\epsilon}>0$. Under the Gaussian approximation, $n-n_{0}e^{\epsilon}\sim\mathcal{N}\left(\lambda\left(1-e^{\epsilon}\right),\lambda\left(1+e^{\epsilon}\right)\right)$. Thus, we can write the probability of triggering a positive event,
$Pr\left(n-n_{0}e^{\epsilon}>0\right)=\frac{1}{2}-\frac{1}{2}\textit{erf}\left(\frac{\lambda\left(e^{\epsilon}-1\right)}{\sqrt{2\lambda\left(1+e^{\epsilon}\right)}}\right),$ (2)
@@
$p_{e}\left(\lambda\right)=\frac{1}{2}-\frac{1}{2}\textit{erf}\left(\frac{\left(\lambda+b_{pr}\right)\left(e^{\epsilon}-1\right)}{\sqrt{2\lambda\left(1+e^{\epsilon}\right)}}\right).$ (3)
@@
Figure 3: a. Displayed static scene. b-c. Synthetic noise event count sampled from the Poisson distribution and generalized negative binomial distribution, respectively. d. Noise event count from a experimental recording. The insets show the histogram for the noise event count.
3.0.1 Sampling Noise Events.
Once we know the probability of observing a single noise event, we can simulate the number of noise events over a fixed time window given the intensity of a scene, $I$. Viewing each potential noise event in this window as a Bernoulli trial with probability $p_{e}$, the noise event count has a binomial distribution with probability of $p_{e}\left(I\right)$ and $N$ trials. With the refractory period (i.e. minimum time between events) much smaller than the inter-event interval, $N$ will be large, and the binomial distribution can be approximated by a Poisson distribution with parameter $N\cdot p_{e}\left(I\right)$.
@@ 4 Reconstructing the Scene
We approach the first challenge by counting the positive and negative polarities of noise events separately. We found that event cameras often set different contrast threshold values to trigger positive and negative events, since the leakage current of the event circuit causes positive events to be triggered more easily [36]. Because of potentially asymmetric behavior of the two polarities, the correlation between light intensity and noise events is also different for each polarity. As shown in Supplementary Fig. 1, the corresponding light intensity becomes more uniquely defined when counting positive and negative events separately.
@@ 4.1 Application on Dynamic Scenes
As in Fig. 1, the static scene reconstruction can also operate in parallel with the event-to-video reconstruction (E2VID) pipeline on recordings with both static and dynamic components, e.g., scenes with moving foreground and static background. We first separate signal events triggered by dynamic changes from other recorded events using an event denoising algorithm [18, 14], and apply an E2VID method to reconstruct the dynamic components. The remaining events can be considered noise events triggered mostly by photon fluctuation, which can be fed into the proposed Noise2Image model for static background reconstruction. For a variable-length recording, we aggregate noise events using a moving window with a fixed temporal width. Lastly, we stitch the dynamic and static scene reconstructions together using a motion mask identified from the signal events.
@@ 5 Experiments
The existing event camera datasets all have a significant amount of scene motion and/or camera motion. It is not possible to fully distinguish between events triggered by intensity changes and noise events [3], and thus they are not suitable for our goal of static scene reconstruction. Hence, we collect our own training/validation dataset, termed noise events-to-image (NE2I). NE2I contains pairs of high-resolution intensity images and noise event recordings from both experimental acquisition and synthetic noise based on the model presented in Sec. 3.0.1.
We image a 24.5 inch LCD monitor displaying static images. In order to calibrate the noise model, we display 256 grayscale values on the monitor and capture both their event response and light intensity. The illuminance of each grayscale value is measured using a light meter placed next to the camera [^1] Since the light meter has optics different from the lens on the event camera, the measured illuminance value may not be equal to but is proportional to the actual illuminance level hitting the camera sensor.. Our event camera is a monochromatic Prophesee Metavision EVK3-HD. The default bias parameters were used for the camera, and no denoising was performed on the raw data. While the event camera has a high pixel count ($1280\times 720$ pixels), it does not have an active pixel sensor (APS) that records frame-based intensity images. To register the event camera recording to the screen, we first imaged a standard checkerboard flashing on the monitor to induce events. The aggregated event count is used to establish a transformation matrix [8], which is applied to any event recordings to align it spatially with the monitor.
@@
Given a sequence of noise events, we first aggregate them into a 2D-matrix of event count. To account for the positive and negative polarities, we separately aggregate them and store as two channels. We train a U-net to map an event count matrix into the corresponding intensity image. A modified U-net from [25] is used for enhanced performance. The network is trained using either experimental training data (described above) or synthetic event counts generated by the statistical noise model in Sec. 3.0.1. For experimental data training, we augment the input noise count by choosing a random starting time for the 1-second window over a 10-second event recording. For synthetic data training, event counts are re-sampled each time on-the-fly.
We use E2VID as a baseline method for the evaluation of static scene recovery even though it was not trained by such data. We compare to the pre-trained model implementation [16] for the original E2VID [42], E2VID+ [46], FireNet [44], FireNet+ [46], ET-Net [53], SPADE-E2VID [10], SSL-E2VID [40], HyperE2VID [17]. A sequence of events within a 1-second window is fed into each method, and the last frame of predicted video (5 predicted frames in total) is used for evaluation.
5.0.1 Metrics.

### v2

# Noise2Image: Noise-Enabled Static Scene Recovery for Event Cameras

@@
Noise2Image: Noise-Enabled Static Scene Recovery for Event Cameras
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Event cameras, also known as dynamic vision sensors, are an emerging modality for measuring fast dynamics asynchronously. Event cameras capture changes of log-intensity over time as a stream of ‘events’ and generally cannot measure intensity itself; hence, they are only used for imaging dynamic scenes. However, fluctuations due to random photon arrival inevitably trigger noise events, even for static scenes. While previous efforts have been focused on filtering out these undesirable noise events to improve signal quality, we find that, in the photon-noise regime, these noise events are correlated with the static scene intensity. We analyze the noise event generation and model its relationship to illuminance. Based on this understanding, we propose a method, called Noise2Image, to leverage the illuminance-dependent noise characteristics to recover the static parts of a scene, which are otherwise invisible to event cameras. We experimentally collect a dataset of noise events on static scenes to train and validate Noise2Image. Our results provide a novel approach for capturing static scenes in event cameras, solely from noise events, without additional hardware.
@@ 1 Introduction
Event cameras [1], also known as neuromorphic cameras or dynamic vision sensors, are an emerging modality for capturing dynamic scenes. Their ability to capture data at a much faster rate than conventional cameras has led to applications in high-speed navigation, augmented reality, and real-time 3D reconstruction [2]. Unlike a conventional CMOS camera that outputs intensity images at fixed time intervals, an event camera detects brightness changes at each pixel asynchronously. When the brightness change at a pixel exceeds some threshold, an event is recorded. The output from an event consists of three elements: a timestamp, the spatial coordinate of the triggered pixel, and a binary polarity (indicating whether it is an increase or decrease of brightness).
@@
While this scheme allows event cameras to operate beyond conventional framerates, it makes them blind to the static components of a scene, which induce no brightness changes over time. This issue is especially prevalent when the camera is not moving and therefore provides no information about the static background. Even though event cameras are not designed to capture an intensity image, it is often needed for downstream applications such as initializing motion tracking algorithms [3]. To mitigate this issue, some event cameras include a conventional frame-based sensor in the pixel circuit to simultaneously image both events and traditional intensity images [4, 5]. Alternatively, a frame-based camera can be installed in parallel, using either a beam splitter [6, 7, 8] or additional view registration [9]. Another approach uses a built an opto-mechanical device that rotates a wedge prism to generate events on static scenes [10]. All of these solutions introduce additional hardware and increase cost, complexity, size, and power consumption, highlighting the challenge of capturing static scenes without adding significant overhead.
Figure 1: Overview of Noise2Image: a. An event pixel circuit compares the change of logarithmic light intensity and produces an event when the change passes the contrast threshold. b-e. To illustrate the illuminance-dependency of noise events on static scenes, we simulate the photon arrival process for low or high brightness scenes. The variance of logarithmic intensity is higher for the low brightness scene, and thus its changes from the reference signal triggering the previous event is also more dramatic, resulting much noise events under low brightness. The dotted black lines in d indicate the contrast thresholds for positive and negative events. f. Noise2Image method only takes noise events to reconstruct the static scene intensity. Noise2Image relies on characterizing the relationship between noise events and using learned priors to resolve ambiguities. For scenes with both static and dynamic components, signal events triggered by intensity changes can be fed into existing event-to-video reconstruction methods.
@@
Our work leverages the fact that, even when the scene is static, event cameras still produce noise events. We focus on low and moderate-brightness regimes (e.g., room light, outdoor sunset) [11, 12, 13] in which the dominant source of noise is photon noise — random fluctuations in the photon arrival process. In contrast, the high-brightness regime (e.g., outdoor daylight) includes leakage noise events [14], which we do not model here. While photon noise is well-studied in the context of frame-based scenes, the events triggered by photon noise are commonly treated as general background noise activity to be filtered out.
@@
In this work we propose a method called Noise2Image that reconstructs a static scene from its event noise statistics, with no hardware modifications. First, we derive a statistical noise model describing how noise event generation correlates with scene intensity, which shows a good correspondence with our experimental measurements. Unlike in conventional sensors, where photon noise grows with the signal, we find that for event cameras, the number of events triggered by photon noise is mostly negatively correlated with the illuminance level due to the logarithmic sensitivity of the sensor. Imaging the static scene then amounts to inverting this intensity-to-noise process. However, the mapping is one-to-many, so not directly invertible; thus, we rely on a learned prior to resolve ambiguities. To train and validate our method, we experimentally collect a dataset of event recordings on static scenes. We demonstrate that Noise2Image can recover photo-realistic images from noise events alone, or can be used to recover the static parts of scenes with dynamics. In addition to testing on in-distribution data, we demonstrate the robustness of this approach with out-of-distribution testing data and live scenes.
@@ 2 Related Work
2.1 Event-to-video reconstruction
Event-to-video reconstruction (E2VID) is a class of methods that recover high frame-rate video from event recordings. These methods can capture high-speed dynamics without motion blur. Because event cameras only capture scene changes, many E2VID approaches include one or two traditional frame-based images, using events to deblur [15, 16, 17, 18], synthesize adjacent frames [19], or interpolate temporally [20, 21, 22, 23].
In contrast, a more challenging class of E2VID reconstruction only uses events for video reconstruction. Because the initial scene intensity is unknown and only the relative intensity changes are measured, reconstruction requires either explicit modeling of spatiotemporal relationships [24] or deep neural networks as a data prior to fill in the missing information [25, 26, 27, 28, 29, 30, 31]. Because it is difficult to collect event recordings paired with the corresponding frame-based videos at scale, the data prior is obtained through synthetic data generation [25, 26] using an event camera simulator [32]. While E2VID methods are effective for recordings with object motion or camera motion, the reconstruction of static scenes is still out of reach since no events should be triggered without motion. Our Noise2Image method can be considered complementary to E2VID; one might use Noise2Image to recover the static parts of the scene and E2VID to recover the dynamic parts.
2.2 Event camera noise characterization
An event recording often contains a number of events not associated with intensity changes, which are termed noise events or background activity [33, 12]. Noise events are attributed to two main sources: photon noise and leakage current [14, 13, 11]. Photon noise is the dominant source of noise in low-brightness conditions [34, 13], while leakage noise events dominate in high-brightness conditions [14, 13]. Although the noise model of event cameras is less studied than traditional sensors, in the lower-light regime, it is generally believed that noise events become less likely to trigger as intensity increases [35, 36]. In event camera simulators, the events triggered by photon noise are modeled as a Poisson process, with the noise event rate linearly decreasing with intensity [36]. The intensity dependency of noise events has been experimentally measured in [13], and shows a non-monotonic relationship with intensity in low light. Our experimental results are consistent with this trend, and we derive a theoretical noise model which explains the relationship between intensity and noise events.
2.3 Event denoising
@@
2.4 Event camera datasets
@@ 3 Modeling Noise Event Statistics
where $t_{0}$ is the timestamp for the most recent event at the same pixel. The logarithmic sensitivity ensures that this triggering condition is adaptive to the brightness level, e.g., a tenfold increase from 1 lux to 10 lux would result in a similar response as a tenfold increase from 100 lux to 1000 lux. As pixels operate asynchronously, we will proceed by considering events at an arbitrary pixel and drop the $\left(x,y\right)$ notation for simplicity.
@@
To trigger a positive event, we need $\log\left(n\right)-\log\left(n_{0}\right)>\epsilon$, which can be re-written as $n-n_{0}e^{\epsilon}>0$. Under the Gaussian approximation, $n-n_{0}e^{\epsilon}\sim\mathcal{N}\left(\lambda\left(1-e^{\epsilon}\right),\lambda\left(1+e^{2\epsilon}\right)\right)$. Thus, we can write the probability of triggering a positive event,
$Pr\left(n-n_{0}e^{\epsilon}>0\right)=\frac{1}{2}-\frac{1}{2}\textit{erf}\left(\frac{\lambda\left(e^{\epsilon}-1\right)}{\sqrt{2\lambda\left(1+e^{2\epsilon}\right)}}\right),$ (2)
@@
$p_{e}\left(\lambda\right)=\frac{1}{2}-\frac{1}{2}\textit{erf}\left(\frac{\left(\lambda+b_{pr}\right)\left(e^{\epsilon}-1\right)}{\sqrt{2\lambda\left(1+e^{2\epsilon}\right)}}\right).$ (3)
@@
Figure 3: a. Displayed static scene. b-c. Synthetic noise event count sampled from the Poisson distribution and generalized negative binomial distribution, respectively. d. Noise event count from a experimental recording. The insets show the histogram for the noise event count. The frequency of the histogram (y-axis) is in log scale.
3.1 Sampling Noise Events
Once we know the probability of observing a single noise event, we can simulate the number of noise events over a fixed time window given the intensity of a scene, $I$. Viewing each potential noise event in this window as a Bernoulli trial with probability $p_{e}$, the noise event count has a binomial distribution with probability of $p_{e}\left(I\right)$ and $N$ trials. With the refractory period (i.e., minimum time between events) much smaller than the inter-event interval (i.e., the time interval between two consecutive noise events at the same pixel), $N$ will be large, and the binomial distribution can be approximated by a Poisson distribution with parameter $N\cdot p_{e}\left(I\right)$.
@@ 4 Reconstructing the Scene
We approach the first challenge by counting the positive and negative polarities of noise events separately. We found that event cameras often set different contrast threshold values to trigger positive and negative events, since the leakage current of the event circuit causes positive events to be triggered more easily [14]. Because of potentially asymmetric behavior of the two polarities, the correlation between light intensity and noise events is also different for each polarity. As shown in Fig. S1, the corresponding light intensity becomes more uniquely defined when counting positive and negative events separately.
@@ 4.1 Application on Dynamic Scenes
As in Fig. 1f, the static scene reconstruction can also operate in parallel with the event-to-video reconstruction (E2VID) pipeline on recordings with both static and dynamic components, e.g., scenes with moving foreground and static background. We first separate signal events triggered by dynamic changes from other recorded events using an event denoising algorithm [44, 35], and apply an E2VID method to reconstruct the dynamic components. The remaining events can be considered noise events triggered mostly by photon fluctuation, which can be fed into the proposed Noise2Image model for static background reconstruction. For a variable-length recording, we aggregate noise events using a moving window with a fixed temporal width. Lastly, we stitch the dynamic and static scene reconstructions together using a binary motion mask identified from the signal events.
@@ 5 Experiments
The existing event camera datasets all have a significant amount of scene motion and/or camera motion. It is not possible to fully distinguish between events triggered by intensity changes and noise events [43], and thus they are not suitable for our goal of static scene reconstruction. Hence, we collect our own training/validation dataset, termed noise events-to-image (NE2I). NE2I contains pairs of high-resolution intensity images and noise event recordings from both experimental acquisition and synthetic noise based on the model presented in Sec. 3.1.
We image a 24.5 inch LCD monitor displaying static images. In order to calibrate the noise model, we display 256 grayscale values on the monitor and capture both their event response and light intensity. The illuminance of each grayscale value is measured using a light meter placed next to the camera. Our event camera is a monochromatic Prophesee Metavision EVK3-HD. The default bias parameters were used for the camera, and no denoising was performed on the raw data. While the event camera has a high pixel count ($1280\times 720$ pixels), it does not have an active pixel sensor (APS) that records frame-based intensity images. To register the event camera recording to the screen, we first imaged a standard checkerboard flashing on the monitor to induce events. The aggregated event count is used to establish a transformation matrix [53], which is applied to any event recordings to align it spatially with the monitor.
@@
Given a sequence of noise events, we first aggregate them into a 2D-matrix of event count. To account for the positive and negative polarities, we separately aggregate them and store as two channels. We train a U-net to map an event count matrix into the corresponding intensity image. A modified U-net from [55] is used for enhanced performance. The network is trained using either experimental training data (described above) or synthetic event counts generated by the statistical noise model in Sec. 3.1. For experimental data training, we augment the input noise count by choosing a random starting time for the 1-second window over a 10-second event recording. For synthetic data training, event counts are re-sampled each time on-the-fly. After training, the model takes around 100ms to recover a static scene on an NVIDIA RTX3090 GPU.
We use E2VID as a baseline method for the evaluation of static scene recovery even though it was not trained with such data. Rather than quantifying our performance gain, this comparison aims to test whether E2VID can identify the correlation between noise events and illuminance level. We compare to the pre-trained model implementation [56] for the original E2VID [25], E2VID+ [26], FireNet [27], FireNet+ [26], ET-Net [28], SPADE-E2VID [29], SSL-E2VID [30], HyperE2VID [31]. A sequence of events within a 1-second window is fed into each method, and the last frame of predicted video (5 predicted frames in total) is used for evaluation.
5.1 Metrics